Circular No. 5701

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993C IN NGC 2954 A. V. Filippenko and T. Matheson, University of California at Berkeley,

report that CCD spectra (range 320-1000 nm) obtained on Jan. 30 UT with the Shane 3-m reflector at Lick Observatory reveal that SN 1993C (cf. IAUC 5699) is of type Ia, probably ∼ 1-2 months past maximum brightness.

PSR J0437–4715 W. Becker, J. Trümper, K. T. S. Brazier, and T. Belloni, Max-Planck-

Institut für Extraterrestrische Physik, communicate: “We report the detection with ROSAT of the 5.75-ms binary pulsar PSR J0437–4715 in the energy band 0.1-2.4 keV. This source was observed from 1990 July 30 to 1990 Aug. 2, for a total exposure of 680 s, as part of the ROSAT PSPC all-sky survey. It is visible as a highly significant point source with a PSPC net countrate of 0.12 ± 0.02 counts/s. The observed flux in the band 0.1-2.4 keV is ∼ 1 × 10^{−12} erg cm^{−2} s^{−1}. Adopting a distance of 140 pc, as determined from the radio dispersion measure, the isotropic x-ray luminosity is 3 × 10^{30} erg/s. The source counts are confined below PSPC channel 40, indicating a rather soft and unabsorbed spectrum. Using the radio frequency extrapolated from the ephemeris given in the Princeton Pulsar Catalogue (cf. also Johnston et al. 1993, to appear in Nature), we detected a pulsed signal at a significance level of 99.96 percent, including the trials involved in optimizing the longitude of periastron to 5^{o}.85. This angle is within the error range reported in the available ephemeris. Inspecting the blue plates of the U.K. Schmidt optical survey, we found a faint object of mag ∼ 21 at α = 4^{h}37^{m}15^{s}.5, δ = −47^{o}15′08′′ (equinox J2000.0), 2′′ from the radio position. This corresponds to an absolute magnitude of ∼ 15, consistent with either a white dwarf or a late-type M star. The positional coincidence makes this object a possible candidate for the binary companion.”

GRO J0422+32 Further visual magnitude estimates (the first two by Merlin correct

those given on IAUC 5690): 1992 Dec. 16.77 UT, 14.2 (J.-C. Merlin, Le Creusot, France); 29.92, 14.2 (Merlin); Jan. 14.81, 14.2 (Merlin); 26.13, 14.0 (J. E. Bortle, Stormville, NY).

1993 February 2 Daniel W. E. Green